aaye e5en as kfh6t 6 1c 3 d , g
Answer #1

Formula sheet

A1 B C D E F G H I J K
2
3 First WACC Needs to be calculated to value the project.
4
5 Formula for WACC is given as:
6 WACC = r(E) × w(E) + r(P) × w(P)+r(D) × (1 – t) × w(D)
7 Where, r(E), r(P) and r(D) are cost of equity, preferred stock and cost of debt, w(E), w(P) and W(D)
8 are weight of equity, preferred stock and debt and t is the tax rate
9
10 Calculation of cost of debt of coupon bonds:
11 Cost of debt will be the yield to maturity of the bond can be calculated as follows:
12 Time to maturity 20 years
13 Annual coupon rate 0.09
14 Par value 1000
15 Market Price 1085
16 Annual coupon =D14*D13
17 Annual Period =D12
18
19 Rate(nper,pmt,PV, [fv],type) function of excel can be used to find the yield to maturity as follows:
20 NPER =D17
21 PMT =D16
22 PV =-D15
23 FV =D14
24
25 Yield to maturity of the coupon bond is =RATE(D20,D21,D22,D23) =RATE(D20,D21,D22,D23)
26
27 Hence Cost of Debt is =D25
28
29
30 Calculation of Cost of Equity:
31
32 Beta of stock 2.45
33 Market Return (rm) 0.12
34 Risk free rate 0.05
35 As Per CAPM
36 r(E) = rf + ?*(rm-rf)
37 Cost of common equity= =D34+D32*D33 =D34+D32*(D33-D34)
38
39 Hence cost of Equity is =D37
40
41
42 Calculation of cost of preferred stock:
43 Cost of preferred stock can be calculated as follows:
44 Annual Dividend of preferred stock =50*5.5% (Assuming $50 as par value)
45 Current Price 40
46 Floatation cost (F) 0
47 Cost of preferred stock =Dividend/Current Price*(1-F)
48 Cost of preferred stock =D44/(D45*(1-D46)) =D44/(D45*(1-D46))
49
50 Hence Cost of Preferred Stock is =D48
51
52
53 Calculation of WACC:
54
55 Source of capital Number of securities Price MV Capital Structure Cost
56 Debt 2000 =D15 =D56*E56 =F56/$F$59 =D27
57 Common Stock 60000 45 =D57*E57 =F57/$F$59 =D39
58 Preferred Stock 36000 =D45 =D58*E58 =F58/$F$59 =D50
59 Total =SUM(F56:F58) =SUM(G56:G58)
60 Tax Rate 0.34
61
62 WACC of the firm is = r(E) × w(E) + r(P) × w(P)+r(D) × (1 – t) × w(D)
63 =G57*H57+G58*H58+G56*H56*(1-G60) =G57*H57+G58*H58+G56*H56*(1-G60)
64
65 Hence WACC of the firm is =D63
66
67 Calculation of NPV:
68 To calculate NPV of the project, free cash flow needs to be calculated as follows:
69 Free Cash Flow = Operating Cash Flow - Capital Expenditures - Change in working capital
70 Operating Cash Flow = EBIT*(1-Tax Rate)+Depreciation
71 Tax Rate 0.34
72 Initial investment calculation:
73 Capital cost required 200000
74
75
76 Depreciation each year can be calculated as follows:
77 Capital cost (B) =D73
78 Depreciation follows MACRS 5 year-half year convention.
79
80 Hence depreciation each year can be calculated as follows:
81 Year 1 Year 2 Year 3 Year 4 Year 5 Year 6
82 MACRS 5 Year depreciation rate (rt) 0.2 0.32 0.192 0.1152 0.1152 0.0576
83 Depreciation (B*rt) =$D77*E82 =$D77*F82 =$D77*G82 =$D77*H82 =$D77*I82 =$D77*J82
84 Book Value =D77 =D84-E83 =E84-F83 =F84-G83 =G84-H83 =H84-I83 =I84-J83
85
86 Net Proceed from sale of machine calculation:
87
88 Proceed from sale of machine at the end of 6th year 50000
89 Book Value of Machine at the end of 6th year =J84
90 Gain or Loss on sale of Systems and Softwares =Proceed From Sale - Book value at the end of sale
91 =D88-D89
92
93 Gain or Loss on sale of Machine =D91
94 Tax on Gain & Loss =D93*D71
95 Net Proceed from Sale of System & Softwares =Proceed from Sale - Tax Expense on gain or loss
96 =D88-D94
97
98 Free cash flow can be calculated as followed:
99 Year 0 1 2 3 4 5 6
100 Investment =-D73
101 Savings 50000 =E101 =F101 =G101 =H101 =I101
102 Depreciation =-E83 =-F83 =-G83 =-H83 =-I83 =-J83
103 Operating Income Before Tax (EBIT) =SUM(E101:E102) =SUM(F101:F102) =SUM(G101:G102) =SUM(H101:H102) =SUM(I101:I102) =SUM(J101:J102)
104 Tax expense =-E103*$D$71 =-F103*$D$71 =-G103*$D$71 =-H103*$D$71 =-I103*$D$71 =-J103*$D$71
105 After Tax operating income (EBIT*(1-T)) =E103+E104 =F103+F104 =G103+G104 =H103+H104 =I103+I104 =J103+J104
106 Add Depreciation =-E102 =-F102 =-G102 =-H102 =-I102 =-J102
107 Operating Cash Flow =E105+E106 =F105+F106 =G105+G106 =H105+H106 =I105+I106 =J105+J106
108 Net Proceed from Sale =D96
109 Free Cash Flow =SUM(D100:D108) =SUM(E107:E108) =SUM(F107:F108) =SUM(G107:G108) =SUM(H107:H108) =SUM(I107:I108) =SUM(J107:J108)
110
111 NPV calculation:
112 NPV of the project is present value of future cash flows discounted at required rate of return less the initial investment.
113 Given the following cash flow and WACC, NPV for the project can be calculated as follows:
114 Year 0 1 2 3 4 5 6
115 Free Cash Flow (FCF) =D109 =E109 =F109 =G109 =H109 =I109 =J109
116 WACC (i) =D65
117 (P/F,i,n) for each year =1/((1+$D116)^E114) =1/((1+$D116)^F114) =1/((1+$D116)^G114) =1/((1+$D116)^H114) =1/((1+$D116)^I114) =1/((1+$D116)^J114)
118 Present Value of cash flows = FCF*(P/F,i,n) =E115*E117 =F115*F117 =G115*G117 =H115*H117 =I115*I117 =J115*J117
119 Present value if future cash flows =SUM(E118:J118) =SUM(E118:J118)
120
121 NPV for Project =Present value fo future cash flows - Initial investment
122 =D119+D115 =D119+D115
123
124 Hence NPV of the Project is =D122
125 (Assuming par value of preferred stock is $50.)
126